OpenStudy (macielynn):

how do you work the problem? y=2x+12 and y=3x+9

OpenStudy (anonymous):

2x+12=3x+9 12-9=3x-2x 3=x x=3

jhonyy9 (jhonyy9):

make equal 2x+12 = 3x+9 subtract fro both sides 2x and will get 12 = x+9 so than subtract from both sides 9 and will get 12-9 = x+9-9 3=x hope is understandably bye